# Inverse Design in Nanophotonics via Representation Learning

### v1

@@ Abstract
Inverse design in nanophotonics, the computational discovery of structures achieving targeted electromagnetic (EM) responses, has become a key tool for recent optical advances. Traditional intuition-driven or iterative optimization methods struggle with the inherently high-dimensional, non-convex design spaces and the substantial computational demands of EM simulations. Recently, machine learning (ML) has emerged to address these bottlenecks effectively. This review frames ML-enhanced inverse design methodologies through the lens of representation learning, classifying them into two categories: output-side and input-side approaches. Output-side methods use ML to learn a representation in the solution space to create a differentiable solver that accelerates optimization. Conversely, input-side techniques employ ML to learn compact, latent-space representations of feasible device geometries, enabling efficient global exploration through generative models. Each strategy presents unique trade-offs in data requirements, generalization capacity, and novel design discovery potentials. Hybrid frameworks that combine physics-based optimization with data-driven representations help escape poor local optima, improve scalability, and facilitate knowledge transfer. We conclude by highlighting open challenges and opportunities, emphasizing complexity management, geometry-independent representations, integration of fabrication constraints, and advancements in multiphysics co-designs.
@@
Inverse Design, Neural Representation, Nanophotonics, Machine Learning, Adjoint Methods, Hybrid Methods
@@ 1 Introduction
Designers typically employ full-wave EM solvers such as the finite-difference time-domain (FDTD) method [22] and the finite element method (FEM) [23]. Even with graphics processing unit (GPU) acceleration [24], large three-dimensional (3D) simulations remain time-consuming and memory-intensive. Solver efficiency and scalability set the practical scope of the design process, as simulation costs often limit the range and resolution of design exploration. This bottleneck is compounded by a broader challenge: neither the optimal device geometry (e.g., the pattern of refractive index) nor optimal physical parameters (such as layer thickness or lattice periodicity) are known a priori: they must therefore be optimized [25]. Traditional manual tuning and heuristic searches are slow and scale poorly in the large design spaces of modern applications [20, 25]. Exhaustive trial-and-error is infeasible; only a small subset of candidate designs can be evaluated within realistic time and memory budgets [26]. This motivates our use of representation learning as a conceptual lens: we classify machine learning (ML)-enhanced inverse-design methods not by the application domain or optimizer, but by where ML is applied. Output-side representation learning trains surrogates or end-to-end networks that approximate the forward or inverse EM mapping and replace a part of the physics solver. Input-side representation learning instead learns a low-dimensional design prior—a latent manifold capturing the salient features of manufacturable, high-performance geometries—thus reshaping the search domain. Viewing the field through this solver-versus-geometry lens complements existing ML-in-photonics surveys [27, 28, 29, 30, 31, 32] and highlights opportunities for hybrid and transferable design pipelines.
@@
Figure 1: Output-side versus input-side representation learning in nanophotonic inverse design. Top panels: two complementary learned representations. Output-side representation (left) models the partial differential equation (PDE) solution or a derived optical property: a differentiable surrogate or physics-informed neural network (PINN) emulates a Maxwell’s equations solver and provides analytic gradients that refine candidates directly in the full design space. Input-side representation (right) models the device geometry itself: a generative model compresses layouts into a low-dimensional latent manifold, and optimization proceeds in that manifold while a PDE solver (the surrogate model) supplies the objective. Bottom panel: schematic of the non-convex nanophotonic design landscape. Two representation-learning frameworks guide the search toward a better optimum (blue star): the output-side surrogate delivers rapid, physics-consistent gradients, and the input-side latent prior confines exploration to geometry regions containing high-performance candidates. Taken separately, each paradigm yields faster convergence, fewer full-wave EM simulations, and lower data requirements; their distinct mechanisms and trade-offs are analyzed in detail throughout this review.
@@
To place the representation-learning framework in context, we first review the physics-driven optimization methods that have shaped nanophotonic inverse design. Foremost are gradient-based techniques especially topology optimization (TO) [25, 33]—which have become the cornerstone of nanophotonic design [34, 35, 36, 21, 37]. They can handle thousands of design variables and generate complex non-intuitive designs that outperform traditional ones [26]. A standard TO loop defines an optical figure of merit (FoM) for the desired optical performance and then iteratively updates the refractive-index distribution to maximize (or minimize) the defined objective [34, 35, 36, 37]. Because a single adjoint simulation provides the full gradient, each iteration scales independently of the parameter count [38]. Recent differentiable-solver frameworks go further: they compute these gradients directly during each iteration, so no extra adjoint simulations are needed [39].
@@
As the complexity of the optimization problem grows (e.g., more degrees of freedom, multi-objective optimization), poor local optima may hinder the optimization process [20]. As a result, the final outcome may depend more strongly on the initial guess, and globally optimal solutions may become elusive. Furthermore, while the gradient calculation itself can be efficient, the iterative nature of these methods, requiring repeated full-wave EM simulations, can lead to prohibitive computational costs and scalability issues, especially for large-scale 3D structures or multi-objective problems. Practical limits, such as structure diameters of approximately $500\lambda$($\lambda$ wavelength), have been observed, with simulations for centimeter-scale devices demanding terabytes of memory and years of computation time [40]. Most of the iterative effort of gradient-based optimization is spent on gradient steps: the simulations and gradient evaluations are only used to determine the next guess in the optimization until a local optimum is reached. Standard gradient-based workflows are memoryless: after reaching a local optimum, they do not leverage the record of the search path that could help a later run. All optimizations have to start from scratch, wasting previous efforts and preventing any meaningful knowledge reuse or transfer learning. Density-based TO also tends to yield a final structure with limited insight into the underlying physical mechanisms or more sophisticated knowledge, like the robustness of the design to fabrication imperfections. Many TO workflows assume fixed physical hyperparameters—such as thicknesses of the different material layers or the period of a periodic structure—from the outset, leaving potentially superior designs unexplored. Shape-based parameterizations can vary some physical parameters and produce simpler layouts [4, 41, 42, 43], but the reduced degrees of freedom may cap ultimate performance [37].
@@
Overall, the evolution of inverse design in nanophotonics points to a broader complexity of modern device requirements that neither pure physics-based methods nor pure data-driven models can tackle on their own. This review aims to parse the landscape of existing methods through the lens of representation learning, articulate the limitations of each, and propose how their hybridization can yield more robust, scalable, and insight-driven solutions to the most demanding light–matter engineering challenges. Previous surveys have already linked ML and photonics. Early work, for example, categorized supervised surrogates and end-to-end inverse pipelines for metamaterials and metasurfaces [27, 28]. Others critically appraise nanophotonic inverse design through the lens of deep NNs and generative models [30, 32]. Complementary work surveys ML-assisted global optimization schemes [31] and physics-informed ML in EM and inverse problems more broadly [29]. Finally, [68] examined synergies between ML and TO.
Although prior surveys give a broad overview, they usually treat forward surrogates and generative models separately from heuristic global searches. Here we fill that gap by classifying methods through the lens of representation learning: ML either targets the output space, using surrogates to emulate solver responses, or targets the input space using generative or latent models to parameterize geometry. This single framework supplies a shared vocabulary for comparing techniques, showing how each choice affects data needs, generalization, and interpretability, and it helps position future work across solver-based (Class A) and learning-based (Class B) design strategies.
@@ 2 Class A: Differentiable Solvers and Surrogate Models (Output-Side Representation Learning)
This integrated framework inherently constrains the network to solutions consistent with fundamental physics. Figures 2(a–c) shown an hPINN implementation [75]. Specifically, two NNs are jointly employed: $u(x;\theta_{u})$, representing the EM field, and $\gamma(x;\theta_{\gamma})$, representing the material distribution (e.g., permittivity). The physics-informed loss function $\mathcal{L}_{F}$ penalizes deviations from Maxwell’s equations and explicitly enforces boundary conditions, e.g., including Dirichlet constraints directly imposed on network outputs and periodicity embedded through sinusoidal input features (see Figure 2(b)). These jointly trained networks simultaneously yield predictions of the EM field distribution (Figure 2(c)) and the corresponding optimized material structure. The hPINN, therefore, learns the solution space defined by Maxwell’s equations, realizing output-side representation learning by directly encoding physical fields and material properties.
Figure 2: Gradient-based co-optimization frameworks combining differentiable EM simulators with PINNs and neural surrogates. (a–c) Hard-constrained PINNs: Two NNs, $\hat{u}(\mathbf{x};\theta_{u})$ and $\hat{\gamma}(\mathbf{x};\theta_{\gamma})$, parameterize EM fields and design variables. Training involves a PDE-informed loss function $\mathcal{L}_{F}$ imposed through automatic differentiation. Dirichlet boundary conditions are enforced in the network outputs, and periodic boundary conditions are embedded via sinusoidal input features. (b) Computational domain showing permittivity design region $\Omega_{2}$ (blue) and perfectly matched layers (PML, hatched). (c) Predicted electric field intensity distribution $|E|^{2}$ resulting from the optimized permittivity $\varepsilon$. (d) Neural-adjoint patch solver: Pillar half-width vectors are transformed into dielectric patches, processed by a convolutional NN (CNN) predicting local EM fields. These fields are stitched together and propagated via the angular-spectrum method. The objective intensity $f=|E(z=F)|^{2}$ is back-propagated using automatic differentiation (PyTorch autograd) to iteratively update the pillar geometries. (e, f) GLOnet + WaveY-Net framework for global TO: (e) GLOnet generates metagrating designs from latent noise vectors; these designs are evaluated by the differentiable WaveY-Net surrogate EM solver. Loss gradients computed by WaveY-Net are back-propagated through GLOnet, enabling differentiable, end-to-end global optimization. (f) WaveY-Net architecture details: A U-Net-based CNN predicts magnetic near-fields, subsequently converted to electric fields via the discrete Ampère’s law. The training loss includes a data-fidelity term ($L_{\text{data}}$) and a Maxwell-residual regularizer ($L_{\text{Maxwell}}$) to keep gradient computations consistent with Maxwell’s equations. (g) Physics-enhanced deep surrogate (PEDS) framework: Fine-resolution geometries are downsampled and combined with coarse-resolution geometries generated by a neural surrogate. The resulting composite structures are evaluated by a fast, low-fidelity solver for rapid performance estimation. High-fidelity solver evaluations of fine-resolution geometries provide offline training data, accelerating the design optimization loop while maintaining physical accuracy. Panels (a–c) reproduced/modified with permission from [75]; panel (d) reproduced/modified with permission from [76]; panels (e, f) reproduced/modified with permission from [59]; panel (g) reproduced/modified with permission from [77].
@@
Extending the representation further, pixel-based differentiable surrogates such as WaveY-Net [59] use PDE losses to directly learn forward solvers from pixel-level geometric inputs. Unlike pillar-based parameterizations, WaveY-Net encodes metasurface geometries explicitly as pixel patterns. It predicts magnetic near-field distributions, then converts them to electric fields through discrete Maxwell relations. The WaveY-Net training incorporates both a data-fidelity term ($\mathcal{L}{\text{data}}$), ensuring accuracy with respect to full-wave simulations and a Maxwell-residual regularizer ($\mathcal{L}{\text{Maxwell}}$) to keep the solution physical. After training, this pixel-based surrogate quickly evaluates device designs, facilitating gradient-based optimization in a fully differentiable manner.
@@
An important enabling factor for many of these output-side methods is the differentiability of the ML components and, increasingly, the entire simulation and optimization pipeline. The neural-adjoint patch solver relies on automatic differentiation for backpropagation (Figure 2(d) [76]). PINNs apply their physics-informed loss via automatic differentiation (Figure 2(a-c) [75]). The GLOnet + WaveY-Net framework achieves fully differentiable end-to-end optimization due to the differentiable nature of the WaveY-Net surrogate (Figure 2(e,f) [59]). The differentiability ensures the flow of gradient information from the final FoM back through the learned output model (and potentially through any generative model proposing the designs) to the design parameters themselves. This capability unlocks gradient-based optimization across complex, ML-augmented systems and signifies a trend towards making increasingly larger portions of the design pipeline differentiable. Output-side representation learning methods have clear advantages. By providing analytical or automatically differentiated gradients (e.g., TO, differentiable solvers) or through rapid surrogate evaluations, these techniques can navigate high-dimensional design spaces more efficiently than gradient-free searches. The learned representation of the solver’s output or its operational behavior is central to this acceleration. They have produced many high-performance photonic devices with complex geometries [21, 34, 35, 36, 37]. Furthermore, the direct incorporation of physical laws in frameworks like PINNs (Figures 2(a-c) [75]) or through physics-based regularization in surrogates like WaveY-Net (Figure 2(f)) [59]) leads to more robust and generalizable learned models of the system’s output. For surrogate models to be truly effective, they must balance computational speed with predictive accuracy. Physics-based regularization helps maintain that fidelity; for instance, WaveY-Net’s Maxwell-residual term ensures its predictions are physically plausible. This helps to mitigate the black-box concerns often associated with purely data-driven ML models.
@@
Despite their successes, direct gradient-based methods and their ML-augmented counterparts face important limitations. They often stall in local optima because they rely on local gradients in a rugged, non-convex design landscape encountered in nanophotonics. Random restarts help, but they add extra computational cost. Many TO workflows fix parameters such as layer thickness or lattice period. This simplifies the search but can skip unconventional, higher-performance designs. Output-side learning, with its focus on the solver, does not inherently resolve this issue of parameterization.
Finally, the limited transfer of knowledge between runs is a drawback. A nanophotonic structure optimized for one set of operating conditions (e.g., a specific wavelength or polarization) may not be easily re-optimized for a different scenario without rerunning the entire optimization loop from scratch. An ML model trained as a surrogate or PINN learns the input-output mapping or the solution to PDEs under the specific assumptions and parameters of its training data. If material properties change significantly, or new physical effects become dominant, the learned model of the output representation may lose its validity. This means that knowledge gained in one design instance is not readily transferable. The computational cost of initially training accurate surrogate models can also be substantial, often requiring a large training or dataset generation costs. These limitations highlight that while output-side representation learning is powerful for accelerating design within a defined physical context, it may not inherently capture more fundamental, transferable design principles, thereby motivating the exploration of input-side representation learning strategies.

### v2

@@ Abstract
Inverse design in nanophotonics, the computational discovery of structures achieving targeted electromagnetic (EM) responses, has become a key tool for recent optical advances. Traditional intuition-driven or iterative optimization methods struggle with the inherently high-dimensional, non-convex design spaces and the substantial computational demands of EM simulations. Recently, machine learning (ML) has emerged to address these bottlenecks effectively. This review frames ML-enhanced inverse design methodologies through the lens of representation learning, classifying them into two categories: output-side and input-side approaches. Output-side methods use ML to learn a representation in the solution space to create a differentiable solver that accelerates optimization. Conversely, input-side techniques employ ML to learn compact, latent-space representations of feasible device geometries, enabling efficient global exploration through generative models. Each strategy presents unique trade-offs in data requirements, generalization capacity, and novel design discovery potentials. Hybrid frameworks that combine physics-based optimization with data-driven representations help escape poor local optima, improve scalability, and facilitate knowledge transfer. We emphasize data efficiency, transferable representations, fabrication-aware design, faster solvers, and hybrid multiphysics co-design.
@@
Inverse Design, Representation Learning, Nanophotonics, Machine Learning, Adjoint Methods, Hybrid Methods
@@ 1 Introduction
Designers typically employ full-wave EM solvers such as the finite-difference time-domain (FDTD) method [22] and the finite element method (FEM) [23]. Even with graphics processing unit (GPU) acceleration [24], large three-dimensional (3D) simulations remain time-consuming and memory-intensive. Solver efficiency and scalability set the practical scope of the design process, as simulation costs often limit the range and resolution of design exploration. This bottleneck is compounded by a broader challenge: neither the optimal device geometry (e.g., the pattern of refractive index) nor optimal physical parameters (such as layer thickness or lattice periodicity) are known a priori: they must therefore be optimized [25]. Traditional manual tuning and heuristic searches are slow and scale poorly in the large design spaces of modern applications [20, 25]. Exhaustive trial-and-error is infeasible; only a small subset of candidate designs can be evaluated within realistic time and memory budgets [26]. This motivates our use of representation learning as a conceptual lens: we classify machine learning (ML)-enhanced inverse-design methods not by the application domain or optimizer, but by where ML is applied. Output-side representation learning trains surrogates or end-to-end networks that approximate the forward or inverse EM mapping and replace a part of the physics solver. Input-side representation learning instead learns a low-dimensional design prior, a latent manifold capturing the salient features of manufacturable high-performance geometries, thus reshaping the search domain. Viewing the field through this solver-versus-geometry lens complements existing ML-in-photonics surveys [27, 28, 29, 30, 31, 32] and highlights opportunities for hybrid and transferable design pipelines.
@@
Figure 1: Output-side versus input-side representation learning in nanophotonic inverse design. Top panels: two complementary learned representations. Output-side representation (left) models the partial-differential-equation (PDE) solution or a derived optical property: a differentiable surrogate or physics-informed neural network (PINN) emulates a Maxwell’s equations solver and provides analytic gradients that refine candidates directly in the full design space. Input-side representation (right) models the device geometry itself: a generative model compresses layouts into a low-dimensional latent manifold, and optimization proceeds in that manifold while a PDE solver (the surrogate model) supplies the objective. Bottom panel: schematic of the non-convex nanophotonic design landscape. Two representation-learning frameworks guide the search toward a better optimum (blue star): the output-side surrogate delivers rapid, physics-consistent gradients, and the input-side latent prior confines exploration to geometry regions containing high-performance candidates. Taken separately, each paradigm yields faster convergence, fewer full-wave EM simulations, and lower data requirements; their distinct mechanisms and trade-offs are analyzed in detail throughout this review.
@@
To place the representation-learning framework in context, we first review the physics-driven optimization methods that have shaped nanophotonic inverse design. Foremost are gradient-based techniques especially topology optimization (TO) [25, 33]—which have become the cornerstone of nanophotonic design [34, 35, 36, 21, 37]. They can handle thousands of design variables and generate complex non-intuitive designs that outperform traditional ones [26]. A standard TO loop defines an optical figure of merit (FoM) for the desired optical performance and then iteratively updates the refractive-index distribution to maximize (or minimize) the defined objective [34, 35, 36, 37]. Because a single adjoint simulation provides the full gradient, each iteration scales independently of the parameter count [38]. Recent differentiable-solver frameworks go further: they compute these gradients directly during each iteration, so no extra adjoint simulations are needed [39, 40].
@@
As the complexity of the optimization problem grows (e.g., more degrees of freedom, multi-objective optimization), poor local optima may hinder the optimization process [20]. As a result, the final outcome may depend more strongly on the initial guess, and globally optimal solutions may become elusive. Furthermore, while the gradient calculation itself can be efficient, the iterative nature of these methods, requiring repeated full-wave EM simulations, can lead to prohibitive computational costs and scalability issues, especially for large-scale 3D structures or multi-objective problems. Practical limits, such as structure diameters of approximately $500\lambda$($\lambda$ wavelength), have been observed, with simulations for centimeter-scale devices demanding terabytes of memory and years of computation time [41]. Most of the iterative effort of gradient-based optimization is spent on gradient steps: the simulations and gradient evaluations are only used to determine the next guess in the optimization until a local optimum is reached. Standard gradient-based workflows are memoryless: after reaching a local optimum, they do not leverage the record of the search path that could help a later run. All optimizations have to start from scratch, wasting previous efforts and preventing any meaningful knowledge reuse or transfer learning. Density-based TO also tends to yield a final structure with limited insight into the underlying physical mechanisms or more sophisticated knowledge, like the robustness of the design to fabrication imperfections. Many TO workflows assume fixed physical hyperparameters, such as thicknesses of the different material layers or the period of a periodic structure, from the outset, leaving potentially superior designs unexplored. Shape-based parameterizations can vary some physical parameters and produce simpler layouts [4, 42, 43, 44], but the reduced degrees of freedom may cap ultimate performance [37].
Beyond these parameterization choices, the additional degrees of freedom in topology optimization come at the cost of fabrication constraints imposed during the design process. For planar devices, manufacturability is typically ensured through projection or filtering schemes that regulate feature sizes and reduce sensitivity to process variations [34]. Recent advances, including foundry-compatible workflows [45], demonstrate that high-performance, fabrication-ready layouts can now be realized within standard lithographic constraints. In free-form three-dimensional architectures, additional limitations arise from structural integrity, which requires coupling electromagnetic optimization with mechanical stability models [46, 47]. Furthermore, extending topology optimization to nonlinear systems introduces new challenges. Although the adjoint method is inherently linear, such cases mainly require deriving a problem-specific adjoint solver [48].
@@
Overall, the evolution of inverse design in nanophotonics points to a complexity that neither pure physics-based methods nor pure data-driven models can tackle alone. Previous surveys have established a strong foundation by linking ML and photonics. Early work, for example, categorized supervised surrogates and end-to-end inverse pipelines for metamaterials and metasurfaces [27, 28]. Others critically appraise nanophotonic inverse design through the lens of deep NNs and generative models [30, 32]. Complementary work surveys ML-assisted global optimization schemes [31] and physics-informed ML in EM and inverse problems more broadly [29]. Finally, ref [72] examined synergies between ML and TO. In essence, these valuable overviews primarily organize the literature either by application domain [27, 28, 26, 25] or by the machine-learning model class employed [30, 32, 31, 29, 72]. This review complements these perspectives by classifying methods through the lens of representation learning: ML either targets the output space, using surrogates to emulate solver responses, or targets the input space, using generative or latent models to parameterize geometry. This solver-versus-geometry approach provides a shared vocabulary for comparing disparate techniques and exposes fundamental trade-offs in data efficiency and generalization. While we present these classes separately for clarity, their power is also realized in hybrid frameworks where input-side generative priors are coupled with output-side surrogate solvers.
@@ 2 Class A: Differentiable Solvers and Surrogate Models (Output-Side Representation Learning)
Figure 2: Gradient-based co-optimization frameworks combining differentiable EM simulators with PINNs and neural surrogates. (a–c) Hard-constrained PINNs: Two NNs, $\hat{u}(\mathbf{x};\theta_{u})$ and $\hat{\gamma}(\mathbf{x};\theta_{\gamma})$, parameterize EM fields and design variables. Training involves a PDE-informed loss function $\mathcal{L}_{F}$ imposed through automatic differentiation. Dirichlet boundary conditions are enforced in the network outputs, and periodic boundary conditions are embedded via sinusoidal input features. (b) Computational domain showing permittivity design region $\Omega_{2}$ (blue) and perfectly matched layers (PML, hatched). (c) Predicted electric field intensity distribution $|E|^{2}$ resulting from the optimized permittivity $\varepsilon$. (d) Neural-adjoint patch solver: Pillar half-width vectors are transformed into dielectric patches, processed by a convolutional NN (CNN) predicting local EM fields. These fields are stitched together and propagated via the angular-spectrum method. The objective intensity $f=|E(z=F)|^{2}$ is back-propagated using automatic differentiation (PyTorch autograd) to iteratively update the pillar geometries. (e, f) GLOnet + WaveY-Net framework for global TO: (e) GLOnet generates metagrating designs from latent noise vectors; these designs are evaluated by the differentiable WaveY-Net surrogate EM solver. Loss gradients computed by WaveY-Net are back-propagated through GLOnet, enabling differentiable, end-to-end global optimization. (f) WaveY-Net architecture details: A U-Net-based CNN predicts magnetic near-fields, subsequently converted to electric fields via the discrete Ampère’s law. The training loss includes a data-fidelity term ($L_{\text{data}}$) and a Maxwell-residual regularizer ($L_{\text{Maxwell}}$) to keep gradient computations consistent with Maxwell’s equations. (g) Physics-enhanced deep surrogate (PEDS) framework: Fine-resolution geometries are downsampled and combined with coarse-resolution geometries generated by a neural surrogate. The resulting composite structures are evaluated by a fast, low-fidelity solver for rapid performance estimation. High-fidelity solver evaluations of fine-resolution geometries provide offline training data, accelerating the design optimization loop while maintaining physical accuracy. Permissions: Panels (a–c) Reproduced with permission.[76] Copyright 2021, Society for Industrial and Applied Mathematics (SIAM). Panel (d) Reproduced under the terms of the CC BY 4.0 license.[77] Copyright 2023, The Author(s). Panels (e, f) Reproduced with permission.[64] Copyright 2022, American Chemical Society. Panel (g) Reproduced with permission.[78] Copyright 2023, Springer Nature.
@@
This integrated framework inherently constrains the network to solutions consistent with fundamental physics. Figures 2(a–c) show an hPINN implementation [76]. Specifically, two NNs are jointly employed: $u(x;\theta_{u})$, representing the EM field, and $\gamma(x;\theta_{\gamma})$, representing the material distribution (e.g., permittivity). The physics-informed loss function $\mathcal{L}_{F}$ penalizes deviations from Maxwell’s equations and explicitly enforces boundary conditions, e.g., including Dirichlet constraints directly imposed on network outputs and periodicity embedded through sinusoidal input features (see Figure 2(b)). These jointly trained networks simultaneously yield predictions of the EM field distribution (Figure 2(c)) and the corresponding optimized material structure. The hPINN, therefore, learns the solution space defined by Maxwell’s equations, realizing output-side representation learning by directly encoding physical fields and material properties.
@@
Extending the representation further, pixel-based differentiable surrogates such as WaveY-Net [64] use PDE losses to directly learn forward solvers from pixel-level geometric inputs. Unlike pillar-based parameterizations, WaveY-Net encodes metasurface geometries explicitly as pixel patterns. It predicts magnetic near-field distributions, then converts them to electric fields through discrete Maxwell relations. The WaveY-Net training incorporates both a data-fidelity term ($\mathcal{L}_{\text{data}}$), ensuring accuracy with respect to full-wave simulations and a Maxwell-residual regularizer ($\mathcal{L}_{\text{Maxwell}}$) to keep the solution physical. After training, this pixel-based surrogate quickly evaluates device designs, facilitating gradient-based optimization in a fully differentiable manner.
@@
An important enabling factor for many of these output-side methods is the differentiability of the ML components and, increasingly, the entire simulation and optimization pipeline. The neural-adjoint patch solver relies on automatic differentiation for backpropagation (Figure 2(d) [77]). PINNs apply their physics-informed loss via automatic differentiation (Figure 2(a-c) [76]). The GLOnet + WaveY-Net framework achieves fully differentiable end-to-end optimization due to the differentiable nature of the WaveY-Net surrogate (Figure 2(e,f) [64]). The differentiability ensures the flow of gradient information from the final FoM back through the learned output model (and potentially through any generative model proposing the designs) to the design parameters themselves. This capability unlocks gradient-based optimization across complex, ML-augmented systems and signifies a trend towards making increasingly larger portions of the design pipeline differentiable. Output-side representation learning methods have clear advantages. By providing analytical or automatically differentiated gradients (e.g., TO, differentiable solvers) or through rapid surrogate evaluations, these techniques can navigate high-dimensional design spaces more efficiently than gradient-free searches. The learned representation of the solver’s output or its operational behavior is central to this acceleration. They have produced many high-performance photonic devices with complex geometries [21, 34, 35, 36, 37]. Furthermore, the direct incorporation of physical laws in frameworks like PINNs (Figures 2(a-c) [76]) or through physics-based regularization in surrogates like WaveY-Net (Figure 2(f)) [64] leads to more robust and generalizable learned models of the system’s output. For surrogate models to be truly effective, they must balance computational speed with predictive accuracy. Physics-based regularization helps maintain that fidelity; for instance, WaveY-Net’s Maxwell-residual term [64] ensures its predictions are physically plausible. This helps to mitigate the black-box concerns often associated with purely data-driven ML models.
@@
Despite their successes, direct gradient-based methods and their ML-augmented counterparts face important limitations. They often stall in local optima because they rely on local gradients in a rugged, non-convex design landscape encountered in nanophotonics. Random restarts help, but they add extra computational cost [21]. Many TO workflows fix parameters such as layer thickness or lattice period. This simplifies the search but can skip unconventional, higher-performance designs [71]. While surrogate models can dramatically accelerate computations, the curse of dimensionality still scales exponentially with the number of design parameters. This acceleration can extend the range of feasible explorations, but asymptotically the exponential growth in data need from the curse of dimensionality dominates, limiting the practical benefit unless the search space is reduced or structured. Moreover, surrogate models are inherently biased toward the distribution of their training data, making knowledge transfer across different material systems, boundary conditions, or target responses difficult. Learning directly from the underlying geometry tends to yield representations that are more transferable than those learned purely from the physics-based outputs, motivating the hybrid and input-side strategies discussed later in this work.